# Protecting Quantum Information via Many-Body Dynamical Localization

### v2

@@ Abstract
Figure 4: (Color online) (a) $\omega_{d}$ with respect to $\Omega$ for the DD of edge spins in the chain of $L=11$. (b) $K_{d}^{(1,1)}$ with respect to $L$. (c) $K$ versus $N_{d}$ for $L=21,41,61$. (d) Evolution of $P_{\uparrow}(j,t)$ in the DD case of $N_{d}=4$ for $K=16\pi/25$ and $L=21$.
Analysis of local DD.— We proceed to present an analysis of the local DD to obtain the required system parameters. For the $j$-th decoupled spin in the MBDL regime, its effective coupling strength is vanishing, corresponding to the local Floquet operator $U^{(j)}_{F}=e^{-i\tilde{K}\sigma_{j}^{z}}e^{-i\Omega T\sigma_{j}^{x}}$, where $\tilde{K}=KT(j-j_{0})^{2}$. When $\tilde{K}=m\pi$ with $m$ being positive integers for certain kick strengths, the operator reduces to $U^{(j)}_{F}=(-1)^{m}e^{-i\Omega T\sigma_{j}^{x}}$. This implies the Rabi oscillation of the effectively decoupled spin with the frequency $\omega_{d}=\Omega T/\pi$. The analytical expression of the oscillation frequency is consistent of the numerical result, as shown in Fig. 4(a). The kick strength required for the DD of $j$-th spin is $K_{d}^{(j,m)}=m\pi/T(j-j_{0})^{2}$. For the centrosymmetric chain of odd spins considered here, the local DD occurs with pairs of spins, such as $j=1,L$ for $L=11$ shown in Fig. 3(d) and Fig. 4. We numerically extract $K_{d}^{(1,1)}$ for various $L$ from the local dynamics and find well agreement with the analytical expression $K_{d}^{(1,1)}=4\pi/T(L-1)^{2}$, as shown in Fig. 4(b).
To store more quantum information without thermalization, one needs more decoupled spins. Remarkably, we obtain the following relation between the kick strength and the DD spin number $N_{d}$ of system size $L$ SM :
@@
In summary, we have explored the disorder-free MBDL in the non-integrable quantum XY chain under periodical kicks. We have revealed the localization phase diagram with dynamical observables and large parameter regions for the MBDL, and uncovered the local DD effect for persistent Rabi oscillation of certain spins. We have proposed the MBDL-protected quantum information at high temperatures, and presented an analysis of the DD to obtain the required system parameters for quantum storage. The MBDL states in spin systems offer a promising platform for studying quantum thermalization and quantum information processing robust against thermal noises.

### v3

@@ Abstract
Figure 4: (Color online) (a) $\omega_{d}$ with respect to $\Omega$ for the DD of edge spins for $L=11$. (b) $K_{d}^{(1,1)}$ with respect to $L$. (c) $K$ versus $N_{d}$ for $L=21,41,61$. Time evolution of $P_{\uparrow}(j,t)$ in the DD case of (d) $L=21,N_{d}=4,K=16\pi/25$, (e) $L=41,N_{d}=8,K=16\pi/25$, and (f) $L=61,N_{d}=8,K=16\pi/49$.
Analysis of local DD.— We proceed to present an analysis of the local DD to obtain the required system parameters. For the $j$-th decoupled spin in the MBDL regime, its effective coupling strength is vanishing, corresponding to the local Floquet operator $U^{(j)}_{F}=e^{-i\tilde{K}\sigma_{j}^{z}}e^{-i\Omega T\sigma_{j}^{x}}$, where $\tilde{K}=KT(j-j_{0})^{2}$. When $\tilde{K}=m\pi$ with $m$ being positive integers for certain kick strengths, the operator reduces to $U^{(j)}_{F}=(-1)^{m}e^{-i\Omega T\sigma_{j}^{x}}$. This implies the Rabi oscillation of the effectively decoupled spin with the frequency $\omega_{d}=\Omega T/\pi$. The analytical expression of the oscillation frequency is consistent of the numerical result, as shown in Fig. 4(a). The kick strength required for the DD of $j$-th spin is $K_{d}^{(j,m)}=m\pi/T(j-j_{0})^{2}$. For the centrosymmetric chain of odd spins considered here, the local DD occurs with pairs of spins, such as $j=1,L$ for $L=11$ shown in Fig. 3(d) and Fig. 4. We numerically extract $K_{d}^{(1,1)}$ for various $L$ from the local dynamics and find well agreement with the analytical expression $K_{d}^{(1,1)}=4\pi/T(L-1)^{2}$, as shown in Fig. 4(b). We further obtain the following relation between the kick strength and the DD spin number $N_{d}$ of system size $L$ SM :
@@
In summary, we have explored the disorder-free MBDL in the non-integrable quantum XY chain under periodical kicks. We have revealed the localization phase regimes with dynamical observables for the MBDL, and uncovered the local DD effect for persistent Rabi oscillation of certain spins. Furthermore, we have proposed the MBDL-protected quantum information at high temperatures, and presented an analysis of the DD to obtain the required system parameters for quantum storage. The MBDL states in spin systems offer a promising platform for studying quantum thermalization and quantum information processing robust against thermal noises.